# Multi-Meta-RAG: Improving RAG for Multi-Hop Queries using Database Filtering with LLM-Extracted Metadata

### v1

@@ Abstract
The retrieval-augmented generation (RAG) enables retrieval of relevant information from an external knowledge source and allows large language models (LLMs) to answer queries over previously unseen document collections. However, it was demonstrated that traditional RAG applications perform poorly in answering multi-hop questions, which require retrieving and reasoning over multiple elements of supporting evidence. We introduce a new method called Multi-Meta-RAG, which uses database filtering with LLM-extracted metadata to improve the RAG selection of the relevant documents from various sources, relevant to the question. While database filtering is specific to a set of questions from a particular domain and format, we found out that Multi-Meta-RAG greatly improves the results on the MultiHop-RAG benchmark. The code is available on GitHub.
@@
large language models retrieval augmented generation multi hop question answering
@@ 1 Introduction
Large Language Models (LLMs) have shown remarkable language understanding and generation abilities [8, 11]. However, there are two main challenges: static knowledge [6] and generative hallucination [3]. Retrieval-augmented generation [4] is an established process for answering user questions over entire datasets. RAG also helps mitigate generative hallucination and provides LLM with a source of new information on which it was not trained [9]. Real-world RAG pipelines often need to retrieve evidence from multiple documents simultaneously, a procedure known as multi-hop querying. Nevertheless, existing RAG applications face challenges in answering multi-hop queries, requiring retrieval and reasoning over numerous pieces of evidence [10]. In this paper, we present Multi-Meta-RAG: an improved RAG using a database filtering approach with LLM-extracted metadata that significantly improves the results on the MultiHop-RAG benchmark.
@@ 2 Related works
In a typical RAG application, we use an external corpus that comprises multiple documents and serves as the knowledge base. Each document within this corpus is segmented into chunks. These chunks are then converted into vector representations using an embedding model and stored in a vector database. Given a user query, RAG typically retrieves the top-K chunks that best match the query. The retrieved chunks, combined with the query are submitted into an LLM to generate a final response.
@@ 3.1 Extraction of Relevant Query Metadata with the LLM
Each question in the MultiHop-RAG [10] benchmark follows a typical structure. Every query requests information from one or more sources of news. In addition, some temporal queries require news articles from a particular date. We can extract the query filter via helper LLM by constructing a few-shot prompt [1] with examples of extracted article sources and publishing dates as a filter. The prompt template is provided in Appendix A. We only run metadata extraction with ChatGPT333gpt-3.5-turbo-1106 because this additional RAG pipeline step must be quick and cheap. We found out that this step takes 0.7 seconds on average for one query.
@@ 4.1 Chunk Retrieval Experiment
We selected two best-performing embedding models from original MultiHop-RAG experiment for testing metadata filtering chunk retrieval performance, bge-large-en-v1.5 [13] and voyage-02 [12]. The retrieved list of chunks is compared with the ground truth evidence associated with each query, excluding the null queries, as they lack corresponding evidence. For evaluation, we assume the Top-K chunks are retrieved and use metrics such as Mean Average Precision at K (MAP@K), Mean Reciprocal Rank at K (MRR@K), and Hit Rate at K (Hit@K). MAP@K measures the average precision of the top-K retrieval across all queries. MRR@K calculates the average of the reciprocal ranks of the first relevant chunk within the top-K retrieved set for each query. Hit@K measures the proportion of evidence that appears in the top-K retrieved set. The experiment (Table 2) with RAG showed considerable improvement in both embeddings for all core metrics MRR@10, MAP@10, Hits@10, and Hits@4. Most notably, for voyage-02 Hits@4 enhanced by 18%. This is important for practical RAG systems, where the top-K retrieved should as as low as possible to account for context window limits and cost.
Table 2: Chunk retrieval experiment results. Top-10 chunks are selected with bge-reranker-large after Top-20 chunks are found via the similarity search and database metadata filtering. A chunk size of 256 and a chunk overlap of 32 is used.
@@
bge-large-en-v1.5 (reported) 0.563 0.4759 0.7183 0.6364
voyage-02 (reported) 0.586 0.4795 0.7467 0.6625
voyage-02 (repository sample) 0.6152 0.2718 a 0.7315 0.6683
@@
Note a
We found a difference between reported MAP@10 and evaluated MAP@10 on the baseline voyage-02 retrieval sample file provided in the MultiHop-RAG repository. We take the evaluated value of 0.2718 for comparisons between RAGs.
@@ 4.2 LLM Response Generation Experiment
Table 3: Generation accuracy of LLMs with Metadata Filtering RAG (Top-6 chunks with voyage-02)
@@
GPT4 (gpt-4-0613) 0.89 0.56 0.63 b c
PaLM (text-bison@001) 0.74 0.47 0.61 b
As with embeddings, we picked two best-achieving LLMs on ground-truth chunks based on MultiHop-RAG initial experiments, GPT-4 and Google PaLM. We achieved substantial improvement in accuracy (Table 3) for both models compared to baseline RAG implementation. Google PaLM accuracy improved by 26% from 0.47 to 0.61. Preliminary GPT-4 results also show a 12% increase from 0.56 to 0.63.
Note b
The accuracy measuring script was not provided in the MultHop-RAG repository. We were able to recreate our script, which shows similar results for ground-truth chunks for both models and on experiments with the sample voyage-02 retrieval list provided in the repository. We accept the answer as correct when the model answer appears in the gold answer or the gold answer appears in the model answer. Both answers are transformed to lowercase, striped from leading and trailing whitespace, and have their punctuation removed before comparison. This script is used for benchmarking Multi-Meta-RAG.
Note c
GPT-4 accuracy is a preliminary result based on 50 queries (2% of all queries). Results on the complete dataset are pending.
@@ 5 Conclusion
In this paper, we introduce Multi-Meta-RAG, a method of improving RAG for multi-hop queries using database filtering with LLM-extracted metadata. Multi-Meta-RAG considerably improves results in both chunk retrieval and LLM generation experiments, while also being quite straightforward and explainable. The proposed solution still has some limitations. Firstly, extracting metadata requires a set of queries from a particular domain and question format, as well as additional inference time. Secondly, it requires the manual creation of a prompt template that will extract the metadata from the query. Thirdly, while the improved results are encouraging, they still fall considerably below the results achieved by feeding LLM precise ground-truth facts. Future work may include testing more generic prompt templates for metadata extraction on a variety of multi-hop datasets from different domains.
Appendix A Metadata Extraction Prompt Template

### v2

@@ Abstract
The retrieval-augmented generation (RAG) enables retrieval of relevant information from an external knowledge source and allows large language models (LLMs) to answer queries over previously unseen document collections. However, it was demonstrated that traditional RAG applications perform poorly in answering multi-hop questions, which require retrieving and reasoning over multiple elements of supporting evidence. We introduce a new method called Multi-Meta-RAG, which uses database filtering with LLM-extracted metadata to improve the RAG selection of the relevant documents from various sources applicable to the question. While database filtering is specific to a set of questions from a particular domain and format, we found that Multi-Meta-RAG greatly improves the results on the MultiHop-RAG benchmark. The code is available on GitHub.
@@
large language models retrieval augmented generation multi-hop question answering
@@ 1 Introduction
Large Language Models (LLMs) have shown remarkable language understanding and generation abilities [10, 13]. However, there are two main challenges: static knowledge [8] and generative hallucination [5]. Retrieval-augmented generation [6] is an established process for answering user questions over entire datasets. RAG also helps mitigate generative hallucination and provides LLM with new information on which it was not trained [11]. Real-world RAG pipelines often need to retrieve evidence from multiple documents simultaneously, a procedure known as multi-hop querying. Nevertheless, existing RAG applications face challenges in answering multi-hop queries, requiring retrieval and reasoning over numerous pieces of evidence [12]. In this paper, we present Multi-Meta-RAG: an improved RAG using a database filtering approach with LLM-extracted metadata that significantly improves the results on the MultiHop-RAG benchmark.
@@ 2 Related works
In a typical RAG application, we use an external corpus that comprises multiple documents and serves as the knowledge base. Each document within this corpus is segmented into chunks. These chunks are then converted into vector representations using an embedding model and stored in a vector database. Given a user query, RAG typically retrieves the top-K chunks that best match the query. The retrieved chunks, combined with the query, are submitted to an LLM to generate a final response.
@@
Several popular benchmarks, such as HotpotQA [16] and 2WikiMultiHopQA [4], can be used for QA from multiple document sources. However, these datasets primarily focus on estimating LLM reasoning skills and do not highlight retrieving evidence from the knowledge base. Another problem is that they are based on Wikipedia, which means LLM’s are already trained on the same data.
Alternative solutions to tackle multi-hop queries include graph-based solutions like Graph RAG [3]. While Graph RAG evaluates MultiHop-RAG dataset, it is used purely as a knowledge base for an independent question set. Another LLM assesses the responses for custom metrics such as comprehensiveness, diversity, empowerment, and directness instead of simple accuracy.
@@ 3.1 Extraction of Relevant Query Metadata with the LLM
Each question in the MultiHop-RAG [12] benchmark follows a typical structure. Every query requests information from one or more sources of news. In addition, some temporal queries require news articles from a particular date. We can extract the query filter via helper LLM by constructing a few-shot prompt [1] with examples of extracted article sources and publishing dates as a filter. The prompt template is provided in Appendix 5.2.1. We only run metadata extraction with ChatGPT333gpt-3.5-turbo-1106 because this additional RAG pipeline step must be quick and cheap. We found out that this step takes 0.7 seconds on average for one query.
@@ 4.1 Chunk Retrieval Experiment
We selected two best-performing embedding models from the original MultiHop-RAG experiment for testing metadata filtering chunk retrieval performance, bge-large-en-v1.5 [15], and voyage-02 [14]. The retrieved list of chunks is compared with the ground truth evidence associated with each query, excluding the null queries, as they lack corresponding evidence. For evaluation, we assume the Top-K chunks are retrieved and use metrics such as Mean Average Precision at K (MAP@K), Mean Reciprocal Rank at K (MRR@K), and Hit Rate at K (Hit@K). MAP@K measures the average precision of the top-K retrieval across all queries. MRR@K calculates the average reciprocal ranks of the first relevant chunk within the top-K retrieved set for each query. Hit@K measures the proportion of evidence that appears in the top-K retrieved set. The experiment (Table 2) with RAG showed considerable improvement in both embeddings for all core metrics: MRR@10, MAP@10, Hits@10, and Hits@4. Most notably, for voyage-02, Hits@4 enhanced by 17.2%. This improvement is important for practical RAG systems, where the top-K retrieved should be as low as possible to account for context window limits and cost.
Table 2: Chunk retrieval experiment results. Top-10 chunks are selected with bge-reranker-large after the top-20 chunks are found via the similarity search and database metadata filtering. A chunk size of 256 and a chunk overlap of 32 is used. We evaluate both Baseline RAG and Multi-Meta-RAG using an evaluation script provided in the MultiHop-RAG repository.
@@
bge-large-en-v1.5 (evaluated) 0.6029 0.2687 0.7490 0.6661
voyage-02 (evaluated) 0.6016 0.2619 0.7419 0.6630
@@ 4.2 LLM Response Generation Experiment
Table 3: Overall generation accuracy of LLMs with MultiHop-RAG (top-6 chunks with voyage-02)
@@
GPT4 (gpt-4-0613) 0.89 0.56 0.606
PaLM (text-bison@001) 0.74 0.47 0.608
As with embeddings, we picked two best-achieving LLMs on ground-truth chunks based on MultiHop-RAG initial experiments, GPT-4 and Google PaLM. We achieved substantial improvement in accuracy (Table 3) for both models compared to baseline RAG implementation. Google PaLM accuracy improved by 25.6% from 0.47 to 0.608. GPT-4 results also show a 7.89% increase from 0.56 to 0.63. The accuracy is calculated by checking if any word in an LLM-generated response is present in the correct gold answer for each question.
Table 4: Generation accuracy of LLMs with MultiHop-RAG per question type (top-6 chunks with voyage-02)
Question Type Accuracy
GPT4 (gpt-4-0613) PaLM (text-bison@001)
Inference 0.951 0.9203
Comparison 0.382 0.5397
Temporal 0.256 0.4545
Null 0.9867 0.2492
Table 4 shows the detailed evaluation results of different question types for GPT-4 and Google PaLM. Both models scored remarkable scores that exceeded 0.9 for inference queries. Google PaLM performs significantly better for comparison and temporal queries than GPT-4. However, PaLM struggles with Null questions, whereas GPT-4 achieves a near-perfect score. These results suggest that combining both models for different queries can be a valid strategy to increase overall accuracy further.
@@ 5 Conclusion
This paper introduces Multi-Meta-RAG, a method of improving RAG for multi-hop queries using database filtering with LLM-extracted metadata. Multi-Meta-RAG considerably improves results in chunk retrieval and LLM generation experiments while being relatively straightforward and explainable compared to alternative solutions, like Graph RAG [3].
5.1 Limitations
The proposed solution still has some limitations. Firstly, extracting metadata requires a set of queries from a particular domain and question format, as well as additional inference time. Secondly, it requires the manual creation of a prompt template that will extract the metadata from the query. Thirdly, while the improved results are encouraging, they still fall considerably below the results achieved by feeding LLM precise ground-truth facts.
5.2 Future work
Future work includes trying more generic prompt templates for metadata extraction using multi-hop datasets from different domains. In addition, testing alternative LLMs, like LLama 3.1 [13], on datasets with more recent cut-off dates is viable.
5.2.1 Acknowledgements
This research was partially funded by OpenAI Researcher Access Program (Application 0000005294).
Appendix
Metadata Extraction Prompt Template